A group of 7 members having a
majority of boys is to be formed out of
6 boys and 4 girls. The number of ways
the group can be formed is
(A) 80 (B) 100 (C) 90 (D) 110

Answer by stanbon(75887)

A group of 7 members having a
majority of boys is to be formed out of
6 boys and 4 girls. The number of ways
the group can be formed is::
-----------------
# of boys must be 4,5,or 6
---
# with 4 boys: [6C4*4C3] = 15*4 = 60
# with 5 boys: [6C5*4C2] = 6*6 = 36
# of 6 boys: [6C6*4C1] = 1*4 = 4
---------------------------------------
# that meet the condition: 60+36+4 = 100
-------------------
Cheers,
Stan H.
===================

(A) 80 (B) 100 (C) 90 (D) 110